OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6).

OpenStudy (anonymous):

Equation of a straight line: y = mx + c Where m is the gradient 2 lines are parallel when the gradient of them both is equal so as long as you have y = -3x then you can put any number after it and the new line would be Parallell e.g. y = -3x + 56434657567099123 y = -3x + 1 For it to pass through ( -1 , 6 ) then put the numbers in and see what you get ( but dont change the gradient ) e.g. y = -3x + 3 ( when x = -1 , y = 6 )

OpenStudy (anonymous):

Well, parallel lines have the same slope (m) and a different y-intercept (b). Slope intercept form is y=mx+b Your original equation is y = -3x + 12. -3 is the slope, and 12 is the y-intercept. To find the y-intercept of the parallel line, we take the coordinates (-1, 6) and plug them into the equation y= -3x+b. It looks like this 6= -3(-1)+b. Solve for b. 6= -3(-1)+b 6= 3+b 3= b So your parallel line equation is y= -3x+3.

OpenStudy (anonymous):

wow @Steph_Rawr352 nice comeback lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

let the equation of required line is m.since this is parallel to the given line hence both has same slope i.e m=-3. hence the equation of required line will be (Y-6)=-3(X-(-1))=>Y=-3X-3+6=>Y=-3X+3...

OpenStudy (anonymous):

LOl thanks to everybody